# Instability of a fluctuating biomimetic membrane driven by an applied uniform DC electric field

### v1

@@ IV.3 Spontaneous curvature
Biomimetic and biological membranes consist of two monolayers which are typically exposed to different environments and may differ in packing density of the lipid molecule’s head and tail [33]. These asymmetry will result into the torque across the membrane and such that the membranes prefer to curve and characterized by the spontaneous curvature. In this study, the asymmetric conductivity of the electrolytes on both sides of the membrane changes the stress distribution, and we will examine the spontaneous curvature $C_{0}$ calculated as the first moment of the lateral stress tensor [18]. After scaling $C_{0}$ with $\lambda_{D}^{-1}$, we will have
[…]
The component $T_{xx}$ can be obtained as
@@
where the base state perssure $p^{(0)}$ in Eq. (26) can also be obtained from $T_{zz}=-p^{(0)}+\frac{1}{2}E_{z}^{(0)2}=0$. Basing on this stress profile, the spontaneous curvature can be obtained as
[…]
If $\gamma=1$, the spontaneous curvature vanishes in a symmetric system, and this corresponds to the result in [18] for the finite-thickness membrane with zero charge density inside. However, the asymmetric conductivities of the bulk fluids tend to curve the flat membrane through the first moment of the stress. $\gamma>1$ corresponds to the conductivity $\sigma_{1}<\sigma_{2}$, and the membrane tends to have a negative curvature. The cell geometry alerted by the asymmetric electric double layers adjacent to the membrane is also reported in experimental work [34].
For current selecting of parameters, with $\beta=10$, $V=1$, $\gamma=2$, $Ca=2$, we will have $\rho_{m}\approx 0.4$, and $|C_{0}|\approx 0.6$. Also, $C_{0}$ increase with $\gamma$ for $\gamma>1$. As $\gamma\rightarrow+\infty$, $|C_{0}|_{\max}=CaV^{2}=2$. The $|C_{0}|\sim O(1)$ indicates that the spontaneous curvature is of the same scale as the inverse Debye length. Thus, in the zero-thickness membrane assumption, we will always have $|C_{0}|d\ll 1$, and it is valid to assume a flat membrane to conduct linear stability analysis.
@@ IV.4 Small perturbation assumption
The infinitesimal perturbation assumption serves as a fundamental principle in the linear stability analysis of membranes or interfaces, a concept employed in EK modeling as seen in our present study and [20], as well as in LDM, as demonstrated in [10]. However, it becomes important to examine the perturbation amplitude relative to the Debye length.
The LDM assumes electroneutrality of the bulk fluid and an infinitesimal amount of free charge confined to the interface. Consequently, the model is inherently constrained by the thin Debye layer limit. Linear analysis derived from the LDM overlooks any dynamic behavior within the Debye layer, implicitly assuming perturbation amplitudes surpassing the Debye layer’s dimensions.
@@
by assuming $h$ to be small. However, when considering the exponential distribution of the base state potential $\phi^{(0)}_{i}$, we encounter $\partial_{z}\phi_{2}^{(0)}=Ae^{\gamma z}$. Substituting $z=h$ into $\partial_{z}\phi_{2}^{(0)}$, we find that to validate the expansion (51), i.e.
@@
the condition $|\gamma h|\ll 1$ is necessary. Similarly, after applying the same procedure to the upper fluid $\partial_{z}\phi_{1}^{(0)}$, we derive the full constraint $|h|\ll\min(\gamma,\gamma^{-1})$ on the perturbation amplitude to ensure the vcalidity of the linear stability analysis. In cases where the conductivity on both sides of the membrane is symmetric, i.e. $\gamma=1$, and $|h|\ll 1$, we intuitively conclude that the linear analysis is only valid for perturbations smaller than the Debye layer.
Now the discrepancy between the linear analyses stemming from the EK and LDM frameworks become apparient. In the EK model, the initial perturbation is assumed to be smaller than the Debye layer ($h\ll 1\ll L/\lambda_{D}$) and intensifies when exposed to instability induced by the electric field near the membrane. Conversely, the LDM presupposes perturbations larger than the Debye layer ($1\ll h\ll L/\lambda_{D}$), with the thin Debye layer riding atop the deformed membrane, which gets stabilized when subjected to a zero electric field in the bulk. This inherent distinction naturally prompts a fundamental question: How can we systematically determine the growth rate that effectively bridges the gap between the EK and LDM approaches, thereby accurately predicting the stability of perturbations with arbitrary amplitudes?” One potential solution is to explore the dynamics beyond the linear regime within the EK framework. By doing so, we anticipate observing the system transitioning to the base state, or other equilibrium states, through different manifolds when subjected to a significant perturbation.
An accompanying question arises: what is the final state of the unstable membrane in the current study, given that the perturbation cannot grow indefinitely due to the conservation of lipid? We anticipate that nonlinearity will drive the system towards another equilibrium state, resulting in the emergence of stationary patterns in the membrane. For instance, [15] predicates a stable state with nonzero curvature as a consequence of the curvature–charge density coupling, where the transition from a flat membrane is induced when the Debye length is increased. This question can also be tackled by exploring nonlinearity, using methods such as fully nonlinear simulations or reduced nonlinear models like amplitude equations or long-wave equations as in [35]. Extending the EK framework to incorporate nonlinear dynamics will be the focus of our future work.
@@ V Conclusion and outlook
The present work investigates the linear instability of a zero-thickness lipid membrane under a DC electric field, subjected to fluids with asymmetric conductivity. An electrokinetic model with the Debye-Hückell approximation is adopted to capture the dynamics of diffuse layers formed near the membrane. The capacitive effect of the membrane is found to be critical for the onset of instability, due to which the charge accumulates inside the Debye layer. Upon perturbation, the charge redistributes and exerts a net force as surface traction on the fluid adjacent to the membrane, leading to system destabilization. While the electrostatic body force tends to stabilize the system, it is unable to counteract the normal traction. The asymmetric conductivity stabilizes the membrane by reducing the base state electric field. This prediction diverges from that of the LDM framework, which overlooks the dynamics of charge within the Debye layer.
@@
The developed analysis is applicable only to perturbations smaller than the Debye layer. It is of interest to develop a model that systematically bridges the regimes of dimensional perturbation $h\ll\lambda_{D}$, $h\sim\lambda_{D}$ and $L\gg h\gg\lambda_{D}$. Such a model would be able to reconcile the contrasting stability predictions made by the present EK framework ($h\sim\lambda_{D}$) and the LDM framework ($L\gg h\gg\lambda_{D}$). One potential approach is to extend the current model into the nonlinear regime, allowing for illustration of the dynamical structure around the base state. Additionally, we anticipate that this nonlinear model will uncover other nontrivial equilibrium states, potentially revealing new patterns within the system.
@@
Another intriguing factor to investigate is the non-equilibrium effect. The present study assumes instantaneous charge relaxation, and under this assumption, the Debye layer remains in equilibrium when subjected to perturbations. However, in dynamic conditions, the bulk conduction process is inherently unsteady and influenced by ion convection due to fluid motion accompanying membrane deformation. In other words, with finite $\alpha$, the perturbation breaks the equilibrium state, causing charge in the bulk to move to or leave the Debye layer. This results in a transient current in the bulk that either charges or discharges the Debye layer. Such transient effects on stability have been studied in previous works [12, 10, 20], where these effects introduce new instabilities in the system. However, in the EK framework, [20] neglects the ion convection and considers a constant current through the bulk and membrane that is independent of the perturbation in the linear order. It would be of interest to include finite $\alpha$ in the current system and investigate how membrane deformation and ion convection modulate instability.
@@ Appendix A Calculation of internal normal traction
since there is no charge inside. The first-order internal potential can be solved as
@@
The $\phi_{m}^{(1)}$ at the membrane can be evaluated through the expansion of continuity of the potential
@@
where external potential and field at the membrane are approximated using values at $z=0$. After substituting Eq. (55) into Eq. (54), the internal normal stress can be obtained as
@@
With $d$ finite, and in the limit $k\rightarrow 0$, we can evaluate

### v2

@@ IV.3 Spontaneous curvature
Biomimetic and biological membranes consist of two monolayers which are typically exposed to different environments and may differ in packing density of the lipid molecule’s head and tail [37]. This asymmetry results in a torque across the membrane and such that the membranes prefer to curve and are characterized by the spontaneous curvature. In this study, the asymmetric conductivity of the electrolytes on both sides of the membrane changes the stress distribution, and we examine the spontaneous curvature $C_{0}$ calculated as the first moment of the lateral stress tensor [18]. After scaling $C_{0}$ with $\lambda_{D}^{-1}$, we have
$C_{0}=Ca\int^{+\infty}_{-\infty}T_{xx}^{(0)}zdz,$ (48)
The component $T_{xx}$ is obtained as
@@
where the base state pressure $p^{(0)}$ in Eq. (26) can also be obtained from $T_{zz}=-p^{(0)}+\frac{1}{2}E_{z}^{(0)2}=0$. Based on this stress profile, the spontaneous curvature is obtained as
$C_{0}=Ca\int^{+\infty}_{-\infty}T_{xx}^{(0)}zdz=\frac{Ca\rho_{m}^{2}}{4}(1-\gamma^{4}).\\ =Ca\beta^{2}V^{2}\frac{1-\gamma^{4}}{\left[\gamma+(1+\gamma^{2})\beta\right]^{2}}.$ (50)
If $\gamma=1$, the spontaneous curvature vanishes in a symmetric system, and this corresponds to the result in [18] for the finite-thickness membrane with zero charge density inside. However, the asymmetric conductivities of the bulk fluids tend to curve the flat membrane through the first moment of the stress. $\gamma>1$ corresponds to the conductivity $\sigma_{1}<\sigma_{2}$, and the membrane tends to have a negative curvature. The cell geometry altered by the asymmetric electric double layers adjacent to the membrane is also reported in experimental work [38].
To illustrate the effect of spontaneous curvature and maintain consistency with earlier figures, we adopt the parameter set $\beta=1$, $V=1$, $\gamma=2$, and $Ca=0.2$. These values fall within the range typically considered in biologically relevant systems. With this choice, $\rho_{m}\approx 0.3$, the resulting spontaneous curvature is $|C_{0}|\approx 0.06$. Since $C_{0}$ increases with $\gamma$ for $\gamma>1$, the value $\gamma=2$ is selected to provide a representative example with non-zero curvature. In the limit $\gamma\to\infty$, the maximum spontaneous curvature reaches $|C_{0}|_{\max}=CaV^{2}=0.2$. The $|C_{0}|\sim O(0.1)$ indicates that the spontaneous curvature is much smaller than the inverse Debye length. Thus, in the zero-thickness membrane assumption, we always have $|C_{0}|d\ll 1$, and it is valid to assume a flat membrane to conduct linear stability analysis [31]. Furthermore, the lateral membrane tension provides a stabilizing force that suppresses out-of-plane deformations. Thus, the influence of spontaneous curvature becomes subleading and does not qualitatively alter the results of the linear stability analysis.
@@ IV.4 Small perturbation assumption
The infinitesimal perturbation assumption serves as a fundamental principle in the linear stability analysis of membranes or interfaces, a concept employed in EK modeling as seen in our present study and Ref. [20], as well as in LDM, as demonstrated in Ref. [10]. However, it becomes important to examine the perturbation amplitude relative to the Debye length.
The LDM assumes electroneutrality of the bulk fluid and free charge confined to the interface. Consequently, the model is inherently constrained by the thin Debye layer limit. Linear analysis derived from the LDM overlooks any dynamic behavior within the Debye layer, implicitly assuming perturbation amplitudes surpassing the Debye layer’s dimensions.
@@
by assuming $h$ to be small. However, when considering the exponential distribution of the base state potential $\phi^{(0)}_{i}$, we encounter $\partial_{z}\phi_{2}^{(0)}=Ae^{\gamma z}$. Substituting $z=h$ into $\partial_{z}\phi_{2}^{(0)}$, we find that to validate Eq. (51), i.e.
@@
the condition $|\gamma h|\ll 1$ is necessary. Similarly, after applying the same procedure to the upper fluid $\partial_{z}\phi_{1}^{(0)}$, we derive the full constraint $|h|\ll\min(\gamma,\gamma^{-1})$ on the perturbation amplitude to ensure the validity of the linear stability analysis. In cases where the conductivity on both sides of the membrane is symmetric, i.e. $\gamma=1$, and $|h|\ll 1$, we intuitively conclude that the linear analysis is only valid for perturbations smaller than the Debye layer.
Now, the discrepancy between the linear analyses stemming from the EK and LDM frameworks become apparent. In the EK model, the initial perturbation is assumed to be smaller than the Debye layer ($h\ll 1\ll L/\lambda_{D}$) and intensifies when exposed to instability induced by the electric field near the membrane. Conversely, the LDM presupposes perturbations larger than the Debye layer ($1\ll h\ll L/\lambda_{D}$), with the thin Debye layer riding atop the deformed membrane, which gets stabilized when subjected to a zero electric field in the bulk. This inherent distinction naturally prompts a fundamental question: How can we systematically determine the growth rate that effectively bridges the gap between the EK and LDM approaches, thereby accurately predicting the stability of perturbations with arbitrary amplitudes?” One potential solution is to explore the dynamics beyond the linear regime within the EK framework. By doing so, we anticipate observing the system transitioning to the base state, or other equilibrium states, through different manifolds when subjected to a significant perturbation.
An accompanying question arises: what is the final state of the unstable membrane in the current study, given that the perturbation cannot grow indefinitely due to the conservation of lipid molecules? We anticipate that nonlinearity will drive the system towards another equilibrium state, resulting in the emergence of stationary patterns in the membrane. For instance, Ref. [15] predicts a stable state with nonzero curvature as a consequence of the curvature–charge density coupling, where the transition from a flat membrane is induced when the Debye length is increased. This question can also be tackled by exploring nonlinearity, using methods such as fully nonlinear simulations or reduced nonlinear models like amplitude equations or long-wave equations as in Ref. [39]. Extending the EK framework to incorporate nonlinear dynamics will be the focus of our future work.
@@ V Conclusion and outlook
The present work investigates the linear instability of a zero-thickness lipid membrane under a DC electric field, separating fluids with asymmetric conductivity. An electrokinetic model with the Debye-Hückel approximation is adopted to capture the dynamics of the diffuse layers formed near the membrane. The capacitive effect of the membrane is found to be critical to the onset of the instability, due to which the charge accumulates inside the Debye layer. Upon perturbation, the charge redistributes and exerts a net force as surface traction on the fluid adjacent to the membrane, leading to destabilization. While the electrostatic body force tends to stabilize the system, it is unable to counteract the normal traction. The asymmetric conductivity stabilizes the membrane by reducing the base state electric field. This prediction diverges from that of the LDM framework, which overlooks the dynamics of charge within the Debye layer.
@@
The developed analysis is applicable only to perturbations with amplitude smaller than the Debye layer. It is of interest to develop a model that systematically bridges the regimes of dimensional perturbation $h\ll\lambda_{D}$, $h\sim\lambda_{D}$ and $L\gg h\gg\lambda_{D}$. Such a model would be able to reconcile the contrasting stability predictions made by the present EK framework ($h\sim\lambda_{D}$) and the LDM framework ($L\gg h\gg\lambda_{D}$). One potential approach is to extend the current model into the nonlinear regime, allowing for illustration of the dynamical structure around the base state. Additionally, we anticipate that this nonlinear model will uncover other nontrivial equilibrium states, potentially revealing new patterns within the system.
@@
Another intriguing factor to investigate is the non-equilibrium effect. The present study assumes instantaneous charge relaxation, and under this assumption, the Debye layer remains in equilibrium. However, in dynamic conditions, the bulk conduction process is inherently unsteady and influenced by ion convection due to fluid motion accompanying membrane deformation. In other words, with finite $\alpha$, the perturbation breaks the equilibrium state, causing charge in the bulk to move to or leave the Debye layer. This results in a transient current in the bulk that either charges or discharges the Debye layer. Such transient effects on stability have been studied in previous works [12, 10, 20], where these effects introduce new instabilities in the system. However, in the EK framework, Ref. [20] neglects the ion convection and considers a constant current through the bulk and membrane that is independent of the perturbation in the linear order. It would be of interest to include finite $\alpha$ in the current system and investigate how membrane deformation and ion convection modulate instability.
This study explores the linear electrohydrodynamic response of a lipid membrane under a DC electric field through a minimal model designed to isolate the essential physical mechanisms. While motivated by the behavior of biological membranes, the model is not intended to reproduce the full complexity of living cells. Rather, it provides a simplified theoretical framework to gain insight into fundamental membrane-scale electrohydrodynamics. In this regard, artificial systems such as giant unilamellar vesicles (GUVs), which permit precise control over experimental conditions, offer a more directly relevant context for comparison [22, 4]. For analytical tractability, the model omits several biophysical factors that may influence membrane behavior. In particular, it neglects in-plane lipid flows and intermonolayer friction, which can introduce dissipation and modulate the membrane’s dynamic response, especially in the nonlinear regime. Future extensions could incorporate a transport equation for lipid density [40] and include the corresponding dissipative forces [41] in the membrane force balance, enabling a more comprehensive description of membrane dynamics under electric fields.
@@ Appendix A Calculation of internal normal traction
since there is no charge inside. The first-order internal potential is solved as
@@
The $\phi_{m}^{(1)}$ at the membrane is evaluated through the expansion of continuity of the potential
@@
where external potential and field at the membrane are approximated using values at $z=0$. After substituting Eq. (55) into Eq. (54), the internal normal stress is obtained as
@@
With $d$ finite, and in the limit $k\rightarrow 0$, we evaluate